### v2

# Strong Data Processing Properties of Rényi-divergences via Pinsker-type Inequalities

Abstract
We investigate strong data processing inequalities (SDPIs) of the Rényi-divergence between two discrete distributions when both distributions are passed through a fixed channel. We provide a condition on the channel for which the DPI holds with equality given two arbitrary distributions in the probability simplex. Motivated by this, we examine the contraction behavior for restricted sets of prior distributions via $f$-divergence inequalities: We provide an alternative proof of the optimal reverse Pinsker’s inequality for Rényi-divergences first shown by Binette [1]. We further present an improved Pinsker’s inequality for Rényi-divergence based on the joint range technique by Harremoës and Vajda [2]. The presented bound is tight whenever the value of the total variation distance is larger than $\nicefrac{{1}}{{\alpha}}$. By framing these inequalities in a cross-channel setting, we arrive at SDPIs that can be adapted to use-case specific restrictions of input distribution and channel. We apply these results to the Rényi local differential privacy amplification through post-processing by channels that satisfy no local differential privacy guarantee.
I Introduction
[…]
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\biggl{[}\biggl{(}\frac{dP}{dQ}\biggr{)}^{\alpha}\biggr{]}$ (1)
in an AWGN channel is derived in [10] via the Hockey-stick divergence $E_{\gamma}$. Recently, Jin et al. [11] show that the usual upper bound on contraction via total variation fails to hold for Rényi-divergences. They however prove that the universal $\chi^{2}$ lower bound contraction holds for general channels. Zamanlooy and Asoodeh [6] show that if a channel satisfies local differential privacy (LDP), SDPIs can be derived in terms of a channels LDP guarantee. These bounds are further applied to mixing times in [12]. Related, it is shown in [13] that the LDP guarantee of a channel is closely related to its $f$-divergence contraction properties.
Information-theoretic tools for assessing privacy in data processing systems have highlighted the crucial role of the Rényi-divergence in this context: On one hand, Rényi differential privacy (RDP) [14] has recently been adopted as a powerful relaxation of differential privacy. Since, it has been shown to be highly useful for, e.g, composition and privacy accounting [15, 16]. On the other hand, information-theoretic measures that aim to quantify privacy in terms of inference limits can often be shown to reduce to formulations involving Rényi-divergences. An early example of this is the usage of Rényi-divergence of order infinity in the framework of quantitative information flow [17, 18]. More recently, the highly related notion of maximal leakage [19] and its variants and generalizations [20, 21, 22, 23, 24] emerged, and solidify the important role of Rényi-divergence for adversarial approaches to privacy. Interestingly, the framework suggested in [25] unifies many of the above mentioned approaches to quantifying privacy, including maximal leakage, RDP and, in a local setting, Rényi local differential privacy (RLDP). On a related note, the recently proposed pointwise adaption of maximal leakage, pointwise maximal leakage (PML) [21] is shown in [26] to enable inferential guarantees that determine the privacy risk of outcomes by evaluating the Rényi-divergence of order $\infty$ between the prior and posterior distribution of the secret.
Rényi-divergence is itself not an $f$-divergence. However, it can be shown that the Rényi-divergence of any order $\alpha$ can be written as a monotone transform of an $f$-divergence parameterized by $\alpha$ [27]. Using this result, standard $f$-divergence inequalities can be applied to Rényi-divergence. Numerous such inequalities are derived in [27] and subsequent works. Perhaps closest to the approach used in this work, such inequalities are used to derive SDPIs in [28].
While the contraction of $f$-divergences–and hence of Rényi-divergences–of LDP channels is studied in works like [6, 29, 30, 31], there are fewer results that consider the contraction of channels that satisfy no LDP guarantee. For such channels, the aforementioned PML framework continues to allow for a (finite) privacy quantification. Since both frameworks rely on randomization for privacy protection, it is reasonable to assume that PML exhibits a similar relationship between contraction and privacy guarantee as observed for LDP.
Contributions: In this work, we derive SDPI-style bounds for Rényi-divergences aimed at information-theoretic privacy problems. We present a condition on the channel under which it does not contract the Rényi-divergence in general in Section III. In Section IV, we present tools for bounding Rényi-divergences in terms of total variation distance: We present an improved (and often tight) Pinsker’s inequality for Rényi-divergences given $\alpha>1$, as well as an alternative proof to the optimal reverse Pinsker’s inequality for Rényi-divergence first shown in [1]. We apply these tools to obtain results on the data processing properties of Rényi-divergence for discrete channels in Section V. The results show that the restriction of the input distributions can strongly improve the ability to obtain meaningful bounds on contraction. Finally, we illustrate the results by providing an example use-case that shows how these bounds enable an improved assessment of the RLDP privacy amplification through post-processing. The privacy properties of the non-RLDP channel in this bound are quantified by its PML. This dependence establishes a connection between the contraction of Rényi-divergence and the PML privacy parameter. Further, it shows how privacy assessment with LDP and PML can be used synergistically.
II Preliminaries
We use uppercase letters to denote random variables, lowercase letters to denote their realizations and caligraphic letters to denote sets. Specifically, we consider the two random variables $X$ and $Y$, where $X$ represents input data into a transition kernel $P_{Y|X}$ (also referred to as channel or mechanism in the following) that induces a random variable $Y$ at its output. Let $P_{XY}$ denote the joint distribution of $X$ and $Y$. Then we use $P_{XY}=P_{Y|X}\times P_{X}$ to imply that $P_{XY}(x,y)=P_{Y|X=x}(y)P_{X}(x)$ and $P_{Y}=P_{Y|X}\circ P_{X}$ to denote the marginalization $P_{Y}(y)=\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)$. We write $\mathcal{P}_{\mathcal{X}}$ for the probability simplex on the set $\mathcal{X}$. For $N\in\mathbb{N}$, we define $[N]\coloneqq\{1,\dots,N\}$ as the set of positive integers up to $N$. Finally, $\log(\cdot)$ denotes the natural logarithm.
II-A Rényi-divergence, $f_{\alpha}$-divergene and $E_{\gamma}$-divergence
For any convex function $f:(0,\infty)\to\mathbb{R}$ such that $f(1)=0$, define the $f$-divergence between two probability measures $P\ll Q$ [32] as
$D_{f}(P||Q)=\mathbb{E}_{Q}\bigg{[}f\bigg{(}\frac{dP}{dQ}\bigg{)}\bigg{]}.$ (2)
The Rényi-divergence [8] $D_{\alpha}(P||Q)$ between $P$ and $Q$ is a monotone transform of the $f_{\alpha}$-divergence [3]:
Definition 1 ($f_{\alpha}$-divergence).
For any $\alpha>0$, the $f_{\alpha}$-divergence is defined as the $f$-divergence $D_{f_{\alpha}}(P||Q)$, where
$f_{\alpha}(x)=\begin{cases}1-x^{\alpha},\text{ if }\alpha<1,\\ x\log x,\text{ if }\alpha=1,\\ x^{\alpha}-1,\text{ if }\alpha>1.\end{cases}$ (3)
The Rényi-divergence of order $\alpha$ is then obtained by
$D_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\biggl{\{}D_{f_{\alpha}}(P||Q)+1\biggr{\}}.$ (4)
We will use the Hockeystick-divergence $E_{\gamma}$ (from here on simply called $E_{\gamma}$-divergence) as a tool for obtaining bounds on the Rényi-divergence, as similarly done in [10, Section 3.2].
Definition 2 ($E_{\gamma}$-divergence, see, e.g., [27]).
Given some parameter $\gamma>0$, the $E_{\gamma}$-divergence between two discrete distributions $P$ and $Q$ is defined as
$E_{\gamma}(P||Q)=\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-\gamma Q(x)|-\frac{1}{2}|1-\gamma|.$ (5)
For twice differentiable $f$, the relation shown in [33],
$D_{f}(P||Q)=\int_{0}^{\infty}E_{\gamma}(P||Q)f^{\prime\prime}(\gamma)d\gamma,$ (6)
can then be used to obtain any such $D_{f}$-divergence. For $\gamma=1$, the $E_{\gamma}$-divergence is equal to the total variation distance between the two distributions, i.e., $E_{1}(P||Q)=\text{TV}(P||Q)$.
II-B Strong Data Processing Inequalities
All $f$-divergences (including the $f_{\alpha}$-divergence)–and hence also the Rényi-divergence–satisfy the data processing inequality (DPI). The DPI states that for any two probability measures $P_{X},Q_{X}$ on $\mathcal{X}$, and some transition kernel $P_{Y|X}$,
$D_{f}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})\leq D_{f}(P_{X}||Q_{X}).$ (7)
We define the contraction coefficient $\eta_{\alpha}(P_{Y|X})$ of the $\alpha$-Rényi-divergence given a channel $P_{Y|X}$ as the maximum factor between the two sides of the above inequality.
$\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P})$ (8)
$=$ $\sup_{Q_{X}\in\mathcal{P}:D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{D_{\alpha}(P_{Y|X}\circ P_{X}||P_{Y|X}\circ Q_{X})}{D_{\alpha}(P_{X}||Q_{X})},$ (9)
$\eta_{\alpha}(P_{Y|X},\mathcal{P})=\sup_{P_{X}\in\mathcal{P}}\eta_{\alpha}(P_{X},P_{Y|X},\mathcal{P}).$ (10)
Here, $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ denotes a subset of the probability simplex on the alphabet $\mathcal{X}$. In this framework, the DPI states that $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})\leq 1$. We are then interested in the conditions for which $\eta_{\alpha}(P_{Y|X},\mathcal{P})<1$, that is, cases in which the inequality in (7) is strict. In this case, we say a strong data processing inequality (SDPI) holds. Similarly to the above, we define $\eta_{\text{TV}}(P_{Y|X},\mathcal{P})$ as the contraction coefficient of the total variation distance given that input distributions are picked from the set $\mathcal{P}$. In the most general case, an SDPI considers $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. However, we will later see that without any restriction, $\eta_{\alpha}(P_{Y|X})=1$ in many cases. We will therefore derive bounds for cases in which $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$. If the uncertainty set $\mathcal{P}$ is clear from context, we omit it for ease of notation.
III Conditions for $\eta_{\alpha}(P_{Y|X})=1$
The following result provides a condition on the channel $P_{Y|X}$ for which there is no contraction, that is, for which no strong data processing inequality holds, in the case $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$.
Proposition 3.
Consider a discrete channel $P_{Y|X}$. Whenever
$\exists x,x^{\prime}\in\mathcal{X}:\textnormal{supp}\big{(}P_{Y|X=x}\big{)}\cap\textnormal{supp}\big{(}P_{Y|X=x^{\prime}}\big{)}=\emptyset,$ (11)
we have $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
Proof:
See Appendix A. ∎
III-A Connection to the Confusion Graph of $P_{Y|X}$
Proposition 3 states that a channel does not contract the Rényi-divergence whenever there exist two input symbols that cause conditional distributions with disjoint support at the channel output. This condition can equivalently be expressed as requiring a channel’s confusion graph to be incomplete [34]: Define the confusion graph $G=(V,W)$ by $V=\mathcal{X}$ and
$W=\{(x_{i},x_{j})\in\mathcal{X}^{2}\mid\exists\,y\in\mathcal{Y}:x_{i}\in\mathcal{X}^{(y)}\wedge x_{j}\in\mathcal{X}^{(y)}\},$ (12)
where $\mathcal{X}^{(y)}$ for each $y$ denotes the positions at which the channel matrix contains non-zero entries, that is,
$\mathcal{X}^{(y)}\coloneqq\{x\in\mathcal{X}:P_{Y|X=x}(y)>0\}.$ (13)
Heuristically, an edge $(x_{i},x_{j})$ between input symbols in the confusion graph indicates that there exists an output symbol $y\in\mathcal{Y}$ that can be caused by both $x_{i}$ or $x_{j}$ at the channel input. That is, a missing edge (compared to the complete version) indicates the existence of a pair of symbols that can be sent over the channel without being confused with each other. The confusion graph–along with its graph-theoretic dual–is extensively used in zero-error information theory, see e.g. [35, 36] for a detailed account of results and methods.
Note that a fully connected confusion graph is not necessarily equivalent to a strictly positive channel matrix. Proposition 3 in a loose sense states that for sufficiently sparse $P_{Y|X}$, Rényi-divergences do not contract in general. The following examples illustrates this condition on the channel matrices.
Example 4 (Block-diagonal matrices).
Any block-diagonal matrix $\text{diag}(B_{1},\dots,B_{m})$ with each $B_{i}\in\mathbb{R}^{k_{i}\times k_{i}},\,k_{i}\in\mathbb{N}$, satisfies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$.
Example 5 ($k$-singular channels).
Inspired by [37], we say a $N\times N$ channel $P_{Y|X}$ is $k$-singular if
$P_{Y|X=x}(y)=\begin{cases}\frac{1}{k},\text{ if }x\in\mathcal{X}^{(y)}\\ 0,\text{ otherwise.}\end{cases}$ (14)
By Example 4, any $k$-singular channel with $k\leq N/2$ satisfies $\eta_{\alpha}(P_{Y|X},\mathcal{P}_{\mathcal{X}})=1$. For illustration, consider the channels
$P=\begin{bmatrix}\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0\\ \nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}\\ \nicefrac{{1}}{{3}}&0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\\ 0&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}&\nicefrac{{1}}{{3}}\end{bmatrix},\quad Q=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ \nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}.\end{bmatrix}$ (15)
Here, $P$ is $3$-singular and $Q$ is $2$-singular. It is easy to see that the assertion in Proposition 3 applies to $Q$, but not to $P$.
IV Pinsker-type Inequalities for Rényi-divergence
The result in Section III shows that in many cases, the contraction coefficient can not be smaller than one whenever $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$. This motivates us to investigate the behavior of $\eta_{\alpha}$ in cases where $\mathcal{P}\subsetneq\mathcal{P}_{\mathcal{X}}$. As we will see in Section VI, this can be useful in cases in which partial knowledge about the input distributions to a channel is available.
[…]
The following theorem bounds the $f_{\alpha}$-divergence as a function of the total variation between two distributions.
Theorem 6 (Reverse Pinkser’s Inequality for $f_{\alpha}$-divergence, special case of [1, Theorem 1]).
Let $\alpha>1$. For any two distributions $P_{X},Q_{X}\in\mathcal{P}$, define
$\Gamma_{\max}\coloneqq\sup_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)},$ (16)
$\Gamma_{\min}\coloneqq\inf_{P_{X},Q_{X}\in\mathcal{P},\,x\in\mathcal{X}}\frac{P_{X}(x)}{Q_{X}(x)}.$ (17)
Then, the maximum $f_{\alpha}$-divergence corresponding to the $\alpha$-Rényi-divergence $D_{\alpha}(P_{X}||Q_{X})$ between the two distributions is upper bounded by
$D_{f_{\alpha}}(P_{X}||Q_{X})\leq\text{TV}(P_{X}||Q_{X})R_{\alpha}(\Gamma_{\max},\Gamma_{\min}),$ (18)
where
$R_{\alpha}(\Gamma_{\max},\Gamma_{\min})=\frac{\Gamma_{\max}^{\alpha}-1}{\Gamma_{\max}-1}-\frac{1-\Gamma_{\min}^{\alpha}}{1-\Gamma_{\min}}.$ (19)
We provide an alternative proof to [1]. Our proof utilizes the $E_{\gamma}$ identity of $f$-divergences.
Proof:
Fix $Q_{X}$ and $P_{X}$ and let $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$. The proof depends on the fact that $E_{\gamma}(P_{X}||Q_{X})$ can only be non-zero on the interval $(\Gamma_{\min},\Gamma_{\max})$. To see this, observe that
$E_{\gamma}(P_{X}||Q_{X})=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|$ (20)
and whenever $P_{X}(x)-\gamma Q_{X}(x)\leq 0$ for all $x\in\mathcal{X}$ or $P_{X}(x)-\gamma Q_{X}(x)\geq 0$ for all $x\in\mathcal{X}$, we have
$E_{\gamma}(P_{X}||Q_{X})$ $=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{|}P_{X}(x)-\gamma Q_{X}(x)\biggr{|}-\frac{1}{2}|1-\gamma|$ (21)
$=\frac{1}{2}\sum_{x\in\mathcal{X}}\biggl{(}\pm P_{X}(x)\mp\gamma Q_{X}(x)\biggr{)}-\frac{1}{2}|1-\gamma|$ (22)
$=\frac{1}{2}|1-\gamma|-\frac{1}{2}|1-\gamma|=0.$ (23)
This happens exactly when $\gamma\leq\min_{x}\frac{P_{X}(x)}{Q_{X}(x)}\eqqcolon\Gamma_{\min}(P_{X},Q_{X})$ or $\gamma\geq\max_{x}\frac{P_{X}(x)}{Q_{X}(x)}\coloneqq\Gamma_{\max}(P_{X},Q_{X})$. Hence, $E_{\gamma}$ is zero everywhere but on the interval $\gamma\in(\Gamma_{\min}(P_{X},Q_{X}),\Gamma_{\max}(P_{X},Q_{X}))\eqqcolon I_{\geq 0}$. From this, we get
$E_{\gamma}(P_{X}||Q_{X})=\begin{cases}0,&\text{ if }\gamma\notin I_{\geq 0},\\ u(\gamma),&\text{ if }\gamma\in I_{\geq 0}.\end{cases}$ (24)
What remains is to bound $u(\gamma)$ to obtain an upper bound in (6). To do this, note that $\gamma\mapsto E_{\gamma}(P||Q)$ is convex and increasing on $[0,1]$ and convex and decreasing on $(1,\infty)$ [10]. Hence, the maximum value is obtained at $\gamma=1$, where we have $E_{1}(Q_{X}||P_{X})=\Delta$. By convexity, we can then upper bound $u(\gamma)$ by the piecewise linear function
$u(\gamma)\leq g(\gamma)\coloneqq\begin{cases}0,&\text{if }\gamma\notin I_{\geq 0},\\ \Delta\frac{\gamma-\Gamma_{\min}(P_{X},Q_{X})}{1-\Gamma_{\min}(P_{X},Q_{X})},&\text{if }\gamma\in I_{\geq 0}\cap(0,1],\\ \Delta\frac{-\gamma+\Gamma_{\max}(P_{X},Q_{X})}{\Gamma_{\max}(P_{X},Q_{X})-1},&\text{if }\gamma\in I_{\geq 0}\cap(1,\infty).\end{cases}$ (25)
The theorem then follows from maximizing $g(\gamma)$ over $P_{X},Q_{X}\in\mathcal{P}$ and solving
[…]
which yields a bound on the $f_{\alpha}$-divergence by (6). ∎
Note that this proof technique can also be used to obtain the more general result in [1] for any $f$-divergence for which $f$ is twice differentiable by replacing $\gamma^{\alpha-2}$ with the general $f^{\prime\prime}(\gamma)$ in (26). Further, if more information about the distributions is known, more specific bounds on $f(\gamma)$ can potentially be used to obtain more specific inequalities in these cases.
Remark 7.
The bound in Theorem 6 recovers the definition of Rényi-divergence of order infinity, that is, it is asymptotically tight with $D_{\infty}(P_{X}||Q_{X})=\log\Gamma_{\max}$. To see this, let again $\Delta\coloneqq\text{TV}(P_{X}||Q_{X})$ and observe that we can plug in the result form Theorem 6 into (4) to obtain
$D_{\alpha}(Q_{X}||P_{X})\leq\frac{1}{\alpha-1}\log\big{(}\Delta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})+1\big{)}.$ (27)
Since $\Gamma_{\min}\leq 1$ and $\Gamma_{\max}\geq 1$, we get the limit as
$D_{\infty}(Q_{X}||P_{X})$ $\leq\lim_{\alpha\to\infty}\frac{1}{\alpha-1}\log\biggl{(}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})\biggr{)}$ (28)
$=\lim_{\alpha\to\infty}\frac{\alpha}{\alpha-1}\log\Gamma_{\max}=\log\Gamma_{\max}.$ (29)
Further, the bound in [27, Theorem 35] can be obtained from Theorem 6 by assuming the lower bound $\Gamma_{\min}=0$.
Next, we present a generalized Pinsker’s inequality for $f_{\alpha}$-divergences of orders $\alpha>1$.
Theorem 8 (Pinsker’s inequality for Rényi-divergences).
For two mutually absolutely continuous probability distributions $P_{X}$ and $Q_{X}$ and any $\alpha>1$, we have
$D_{f_{\alpha}}(P_{X}||Q_{X})\geq g_{\alpha}\biggl{(}\text{TV}(P_{X}||Q_{X})\biggr{)}$ (30)
where
$g_{\alpha}(t)=\begin{cases}e^{2(\alpha-1)t^{2}}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha<2,\\ (4t^{2}+1)^{\alpha-1}-1,&\text{ if }t<\nicefrac{{1}}{{\alpha}}\text{ and }\alpha\geq 2,\\ (1-t)^{1-\alpha}-1,&\text{ otherwise }\end{cases}$ (31)
Proof:
The proof follows from bounding the joint range [2] of the two divergences and is presented in Appendix B. ∎
Remark 9.
For any $\alpha>1$, the bound in Theorem 8 is realized by the pair of distributions $P_{X}=[0,1]$, $Q_{X}=[t,1-t]$ in the case $t\geq\nicefrac{{1}}{{\alpha}}$. Hence, it is tight for this range of values. It is also tight for all $t\geq 0$ given that $\alpha=2$. In this case, it reduces to the joint-range-based bound between $\chi^{2}$-divergence and total variation distance, as e.g. presented in [3]. As a result of the former, it is also asymptotically tight for $\alpha\to\infty$, since in this case, the condition $t\geq\nicefrac{{1}}{{\alpha}}$ hold for all $t>0$.
Remark 10.
The function $g_{\alpha}$ is injective, but not bijective. That is, in order to obtain an inverse bound (later used in Corollary 11) we can define the piecewise function
$g^{-1}_{\alpha}(s)=\begin{cases}\sqrt{\frac{\log(s+1)}{2(\alpha-1)}},&\text{if}\alpha<2,s<h_{1}(\alpha),\\ \frac{1}{2}\sqrt{(s+1)^{(\frac{1}{\alpha-1})}-1},&\text{if }\alpha\geq 2,s<h_{2}(\alpha)\\ \max\{1-(s+1)^{\frac{1}{1-\alpha}},\nicefrac{{1}}{{\alpha}}\},&\text{otherwise},\end{cases}$ (32)
where $h_{2}(\alpha)\coloneqq(1+4\alpha^{-2})^{\alpha-1}-1$ and $h_{1}(\alpha)\coloneqq 2-\nicefrac{{2}}{{\alpha}}$.
An example of the the presented bound on the joint range for $\alpha=4$ is shown in Figure 1. Other values of $\alpha$ yield a visually similar result, with the discontinuity at $\nicefrac{{1}}{{\alpha}}$ moving closer to the origin as the value of $\alpha$ increases.
Figure 1: Joint range of $D_{f_{\alpha}}(P||Q)$ and $\text{TV}(P||Q)$ for the case $\alpha=4$ as well as the bound presented in Theorem 8.
V Strong Data Processing with Pinsker
[…]
$\Gamma_{\max}=\sup_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (33)
$\Gamma_{\min}=\inf_{P_{X},Q_{X}\in\mathcal{P},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)Q_{X}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X}(x)}$ (34)
[…]
Corollary 11.
For a discrete channel $P_{Y|X}$, a set $\mathcal{P}\subseteq\mathcal{P}_{\mathcal{X}}$ and some $\alpha>1$, let $\eta\coloneqq\eta_{TV}(P_{Y|X})$. We have
$D_{f_{\alpha}}(P_{Y}||Q_{Y})\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(D_{f_{\alpha}}(P_{X}||Q_{X})).$ (35)
Proof:
Follows from Theorem 6 with $P_{Y}=P_{X}\circ P_{Y|X}$ and $Q_{Y}=Q_{X}\circ P_{Y|X}$ and the inequality in Theorem 8. ∎
This result stresses the relation between $\Gamma_{\max}$ and $\Gamma_{\min}$ as formulated in (33) and (34) and the contraction behavior of a channel: If $\Gamma_{\max}-\Gamma_{\min}$ is small, the channel significantly contracts the $f_{\alpha}$-divergence. If the difference is large, contraction will be small. What remains is to bound $\Gamma_{\max}$ and $\Gamma_{\min}$ such that an operational bound is obtained. Such bounds should be use-case specific, as the structure of the uncertainty set $\mathcal{P}$, as well as specifics about the channel are crucial factors. We present one such use-case specific bound in the following.
VI Application: Post-Processing for RLDP
In this section, we show how the above results can be used to obtain improved post-processing inequalities for Rényi local differential privacy (RLDP) [14, 25]. Consider the Markov chain $W-X-Y$, where $W$ is a random variable on a set $\mathcal{W}$ that represents some private data, and the transition kernel $P_{X|W}$ is a privacy mechanism satisfying $(\varepsilon,\alpha)$-RLDP, i.e.,
$\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}D_{\alpha}(P_{X|W=w}||P_{X|W=w^{\prime}})\leq\varepsilon.$ (36)
Further, assume that the transition kernel $P_{Y|X}$ does not satisfy any finite RLDP guarantee, that is, the channel matrix contains zero elements. From the post-processing property of RLDP, it is clear that the privacy guarantee of the system $P_{Y|W}$ will remain bounded by the guarantee of $P_{X|W}$. However, it is reasonable to assume that there exist channels $P_{Y|X}$ containing zero elements that will result in an improved guarantee of the composed system $P_{Y|W}$ if the output of the RLDP $P_{X|W}$ is applied to their input. In fact, using the results derived above, we can bound the $f_{\alpha}$-divergence of $P_{Y|W}$ by
$D_{f_{\alpha}}(P_{Y|W=w}||P_{Y|W=w^{\prime}})$ (37)
$=D_{f_{\alpha}}(P_{Y|X}\circ P_{X|W=w}||P_{Y|X}\circ P_{X|W=w^{\prime}})$ (38)
$\leq\eta R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}\biggl{(}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})\biggr{)},$ (39)
where $\eta\coloneqq\eta_{TV}(P_{Y|X},\mathcal{P}_{\mathcal{X}})$ is again a shorthand for the contraction coefficient of the total variation distance. Since $P_{X|W}$ satisfies $(\varepsilon,\alpha)$-RLDP, we can therefore bound the RLDP guarantee of the entire system by applying (4) as
$\varepsilon^{*}\coloneqq\text{RLDP}(P_{Y|W})\leq\varphi(\varepsilon_{f},\alpha),$ (40)
where $\varepsilon_{f}\coloneqq\max_{w,w^{\prime}}D_{f_{\alpha}}(P_{X|W=w}||P_{X|W=w^{\prime}})$ is the $f_{\alpha}$-divergence corresponding to the $\varepsilon$-valued Rényi-divergence,
$\varphi(\varepsilon_{f},\alpha)\coloneqq\frac{1}{\alpha-1}\log\biggl{(}\eta_{TV}R_{\alpha}(\Gamma_{\max},\Gamma_{\min})g_{\alpha}^{-1}(\varepsilon_{f})+1\biggr{)}.$ (41)
Further, in this setup, the set of input distributions $\mathcal{P}$ is just the set of rows of the mechanism $P_{X|W}$, and each of these distributions is induced by one of the outcomes $W=w$. Hence, the quantities $\Gamma_{\max}$ and $\Gamma_{\min}$ reduce to
$\Gamma_{\max}=\max_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)},$ (42)
$\Gamma_{\min}=\min_{w,w^{\prime}\in\mathcal{W},\,y\in\mathcal{Y}}\frac{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}.$ (43)
While these quantities can be computed directly if both channels are perfectly known, it is useful to obtain a bound that only depends on the privacy guarantees of the channels. The following proposition provides such a bound, where the privacy of the second, non-RLDP channel is quantified by its pointwise maximal leakage (PML) [21]. The proposition shows one way in which the increased flexibility of leakage assessment with PML can complement (L)DP guarantees.
Proposition 12.
If $P_{X|W}$ is a $N\times N$ randomized response mechanisms satisfying $\varepsilon$-LDP according to [31], then
$\Gamma_{\max}\leq\biggl{(}1-\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (44)
$\Gamma_{\min}\geq\biggl{(}1+\frac{e^{\varepsilon}-1}{N+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1},$ (45)
where $\ell_{w}(X\to y)$ denotes the PML from $X$ to an outcome $Y=y$ under the input distribution induced by $W=w$, i.e.,
$\ell_{w}(X\to y)\coloneqq\log\frac{\max_{x\in\mathcal{X}}P_{Y|X=x}(y)}{\sum_{x\in\mathcal{X}}P_{Y|X=x}(y)P_{X|W=w}(x)}.$ (46)
Further, equality is achieved in both bounds if $P_{Y|X}$ has a zero entry in each column, i.e., if $\min_{x}P_{Y|X=x}(y)=0$ $\forall y\in\mathcal{Y}$.
Proof:
See Appendix C. ∎
The following example illustrates this bound.
Figure 2: RLDP and corresponding bound in Example 13.
Example 13.
Let $P_{X|W}$ be a $5\times 5$ randomized response mechanism in [31] satisfying $\log 6$-LDP, and $P_{Y|X}$ given by
$P_{Y|X}=\begin{bmatrix}\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0&0\\ 0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0&0\\ 0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}&0\\ 0&0&0&\nicefrac{{1}}{{2}}&\nicefrac{{1}}{{2}}\\ \nicefrac{{1}}{{2}}&0&0&0&\nicefrac{{1}}{{2}}\end{bmatrix}.$ (47)
[…]
Appendix A Proof of Proposition 3
For $\mathcal{P}=\mathcal{P}_{\mathcal{X}}$, we have
$\eta_{\alpha}(P_{Y|X})=$ (48)
$\sup_{P_{X},Q_{X}\in\mathcal{P}_{X}:\,D_{\alpha}(P_{X}||Q_{X})\neq 0}\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}$ (49)
and the statement follows by constructing $P_{X}$ and $Q_{X}$ with full support such that in the limiting case to the support set boundaries, we get
$\frac{\log\biggl{\{}\sum_{y\in\mathcal{Y}}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)P_{X}(x)\biggr{)}^{\alpha}\biggl{(}\sum_{x\in\mathcal{X}^{(y)}}P_{Y|X=x}(y)Q_{X}(x)\biggr{)}^{1-\alpha}\biggr{\}}}{\log\biggl{\{}\sum_{x\in\mathcal{X}}P_{X}(x)^{\alpha}Q_{X}(x)^{1-\alpha}\biggr{\}}}=1.$ (50)
Since the DPI for Rényi-divergences states that $\eta_{\alpha}(P_{Y|X})\leq 1$ for all $P_{Y|X}$, constructing this distribution proves $\eta_{\alpha}(P_{Y|X})=1$.
Construction of $P_{X}$ and $Q_{X}$: For ease of notation, let $\mathcal{X}\coloneqq[N]$. Further, define the indicator function by
$\delta_{i}(x)=\begin{cases}1,&\text{if }x=i\\ 0,&\text{otherwise}.\end{cases}$ (51)
For some very small $\gamma>0$, let $P_{X}^{\gamma}$ be a perturbed single point mass at $x=i$ for some $i\in[N]$, that is,
$P_{X}^{\gamma}(x)=(1-\gamma)\delta_{i}(x)+\frac{\gamma}{N-1}(1-\delta_{i}(x)),$ (52)
and let
$Q_{X}^{\gamma}(x)=\frac{1-\gamma}{2}(\delta_{i}(x)+\delta_{j}(x))+\frac{\gamma}{N-2}(1-\delta_{i}(x)-\delta_{j}(x)),$ (53)
where $j$ is such that any $\mathcal{X}^{(y)}\coloneqq\{x\in[N]:P_{Y|X=x}(y)>0\}$ that contains $i$ does not contain $j$. Then we get
$\lim_{\gamma\to 0}\eta_{\alpha}(P^{\gamma}_{X},P_{Y|X})$ $=\frac{\log\biggl{\{}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}\biggl{(}P_{Y|X=i}(y)Q^{0}_{X}(i)\biggr{)}^{1-\alpha}\biggl{(}P_{Y|X=i}(y)\biggr{)}^{\alpha}\biggr{\}}}{\log\biggl{\{}Q_{X}^{0}(i)^{1-\alpha}\biggr{\}}}$ (54)
$=\frac{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\sum_{y\in\textnormal{supp}(P_{Y|X=i})}P_{Y|X=i}(y)\biggr{\}}}{\log\biggl{\{}Q^{0}_{X}(i)^{1-\alpha}\biggr{\}}}$ (55)
$=\frac{(1-\alpha)\log Q^{0}_{X}(i)}{(1-\alpha)\log Q^{0}_{X}(i)}=1,$ (56)
where the second to last equality follows from the row-stochasticity of $P_{Y|X}$.
By this construction we have $\eta_{\alpha}(P_{Y|X})\geq 1$. Hence, the data processing inequality for Rényi-divergences implies $\eta_{\alpha}(P_{Y|X})=1$. Noticing that choosing $i$ and $j$ in this manner is only possible if there exists some $i,j\in[N]$ such that $\textnormal{supp}(P_{Y|X=i})\cap\textnormal{supp}(P_{Y|X=j})=\emptyset$ finishes the proof. ∎
Appendix B Proof of Theorem 8
From [27] and the fact that $D_{f_{\alpha}}$ is a monotone transform of the Rényi-divergence of the same order, we know that
$\inf_{P,Q\in\mathcal{P}_{X}:\,TV(P||Q)=t}D_{f_{\alpha}}(P||Q)=\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q),$ (57)
where $d_{f_{\alpha}}(p||q)$ denotes the binary $f_{\alpha}$-divergence
$d_{f_{\alpha}}(p||q)\coloneqq\frac{p^{\alpha}}{q^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-q)^{\alpha-1}}-1.$ (58)
From the definition of the total variation, we get
$\min_{p,q:\,|p-q|=t}d_{f_{\alpha}}(p||q)=\min_{p\in[0,1-t]}\biggl{\{}\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}\biggr{\}}-1.$ (59)
Define
$\psi_{\alpha}(p,t)\coloneqq\frac{p^{\alpha}}{(p+t)^{\alpha-1}}+\frac{(1-p)^{\alpha}}{(1-p-t)^{\alpha-1}}-1.$ (60)
Taking derivatives with respect to $p$ yields
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}$ (61)
and for any $\alpha>1$ and all $t<1,\,p\leq 1-t$,
$\frac{\partial^{2}}{\partial p^{2}}\psi_{\alpha}(p,t)=(\alpha-1)\alpha t^{2}\biggl{[}\frac{p^{\alpha-2}}{(p+t)^{\alpha+1}}+\frac{(1-p)^{\alpha-2}}{(1-p-t)^{\alpha+1}}\biggr{]}\geq 0,$ (62)
from which we can conclude that $\psi_{\alpha}(p,t)$ is convex in $p$.
First, consider the case $t\geq\nicefrac{{1}}{{\alpha}}$. We can bound the first derivative by
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)\geq(\alpha t+p-1)\biggl{(}\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{)}\geq 0,$ (63)
with equality for $t=\nicefrac{{1}}{{\alpha}}$ if $p=0$. Otherwise strict inequality holds. Together with the convexity of $\psi_{\alpha}$ in $p$, this implies that the maximum value of $\psi_{\alpha}$ for all $t\geq\nicefrac{{1}}{{\alpha}}$ will be achieved at $p=0$. At this point, we have
$\psi_{\alpha}(0,t)=(1-t)^{1-\alpha}.$ (64)
Secondly, for $t<\nicefrac{{1}}{{\alpha}}$, we first consider $\alpha\geq 2$. Let $P,Q$ be binary probability distributions with $P(0)=p$ and $Q(0)=p+t$ Then, noticing that $x\mapsto x^{\alpha-1}$ is convex for $\alpha\geq 2$, we apply Jensen’s inequality to obtain
$\psi_{\alpha}(p,t)$ $=\mathbb{E}_{X\sim P}\biggl{[}\biggl{(}\frac{P(X)}{Q(X)}\biggr{)}^{\alpha-1}-1\biggr{]}$
$\geq\biggl{(}\mathbb{E}_{X\sim P}\biggl{[}\frac{P(X)}{Q(X)}\biggr{]}\biggr{)}^{\alpha-1}-1$
$=\biggl{(}1+\frac{t^{2}}{(p+t)(1-p-t)}\biggr{)}^{\alpha-1}-1$
$\geq(1+4t^{2})^{\alpha-1}-1.$
Finally, we bound the remaining case with the standard Pinsker’s inequality: Since the Rényi-divergence is increasing in $\alpha$ [9], so is $D_{f_{\alpha}}$. Hence the standard Pinsker’s inequlity for the Kullback-Leibler divergence $\text{KL}(P||Q)$ (Rényi-divergence of order $\alpha=1$)
$\text{KL}(P||Q)=D_{1}(P||Q)\geq 2\text{TV}(P||Q)^{2}$ (65)
will continue to hold for $\alpha>1$. Utilizing (4), this yields
$D_{f_{\alpha}}(P||Q)\geq\exp\big{(}2(\alpha-1)\text{TV}(P||Q)^{2}\big{)}-1.$ (66)
Putting the derived bounds together to form $g_{\alpha}(t)$ finishes the proof. ∎
Note a tight bound in the case $t<\nicefrac{{1}}{{\alpha}}$ can be obtained by solving
$\frac{\partial}{\partial p}\psi_{\alpha}(p,t)=(\alpha t+p)\biggl{[}\frac{p^{\alpha-1}}{(p+t)^{\alpha}}+\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}\biggr{]}-\frac{(1-p)^{\alpha-1}}{(1-p-t)^{\alpha}}=0.$ (67)
In the range $t\in(0,\nicefrac{{1}}{{\alpha}})$, this equation has a real solution $p^{*}$ that satisfies $p^{*}\in[0,1]$. Due to the convexity of $\psi_{\alpha}$ in $p$, a (tight) lower bound is then given by
$g_{\alpha}(t)=\psi_{\alpha}(p^{*},t).$ (68)
However, it remains unclear if $p^{*}$ can be expressed in closed form. Due to this complication, the result above utilizes more rudimentary bounds in this case.
Appendix C Proof of Proposition 12
Note that for a randomized response mechanism according to [31], we can write for any two rows $w$, $w^{\prime}$,
$P_{X|W=w}(x)=P_{X|W=w^{\prime}}(x)+\kappa(x),$ (69)
where
$\kappa(x)\in\{0,\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1},\frac{1-e^{\varepsilon}}{|\mathcal{X}|+e^{\varepsilon}-1}\}$ (70)
and each non-zero element is taken exactly once. Therefore,
$||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}.$ (71)
Using this observation, we can write
$\Gamma_{\max}$ $=\max_{w,w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w}(x)}$ (72)
$=\max_{\kappa:\,||\kappa||_{1}=2\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}-1}}\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)+\sum_{x}P_{Y|X=x}(y)\kappa(x)}$ (73)
$=\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\biggl{(}\max_{x}P_{Y|X=x}(y)-\min_{x}P_{Y|X=x}(y)\biggr{)}}$ (74)
$\leq\max_{w^{\prime}}\,\max_{y}\frac{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{\prime}}(x)-\frac{||\kappa||_{1}}{2}\max_{x}P_{Y|X=x}(y)}.$ (75)
Now, fix $y$ and let $w^{*}\in\mathcal{W}$ be the solution to the maximization over $\mathcal{W}$. By letting
$\ell_{w^{*}}(X\to y)=\log\frac{\max_{x}P_{Y|X=x}(y)}{\sum_{x}P_{Y|X=x}(y)P_{X|W=w^{*}}(x)},$ (76)
we can further simplify the above bound as
$\Gamma_{\max}\leq\max_{y}\frac{1}{1-\frac{||\kappa||_{1}}{2}\exp\ell_{w^{*}}(X\to y)}=\biggl{(}1-\frac{e^{\varepsilon}-1}{|\mathcal{X}|+e^{\varepsilon}+1}\exp{\max_{y\in\mathcal{Y}}\min_{w\in\mathcal{W}}}\,\ell_{w}(X\to y)\biggr{)}^{-1}.$ (77)
The proof for $\Gamma_{\min}$ follows analogously. ∎